# Secure Smart Contract with Control Flow Integrity

### v1

@@ 6.3. RQ3: Control Flows Introduced by Different Actors
A key insight from this analysis is that a significant number of protocols ($20$, as highlighted in gray) exhibit a relatively low number of non-trivial control flows ($\leq 9$) throughout their operational lifetimes prior to being hacked. Notably, in $10$ protocols (Warp, CheeseBank, ValueDeFi, Eminence, IndexFi, RevestFi, Punk, BlueberryProtocol, PikeFinance, and UwULend), the hack was the first external non-trivial control flow introduced, beyond those generated internally (as indicated by gray cells showing 0 nCF from O-Tx but exactly 1 nCF from E-Tx). This insight suggests that $11$ hacks could be simply prevented with zero false positive rates by restricting ALL external (“non-trusted”) developers, allowing only the protocol deployers to create new control flows.
The analysis reveals that E-Tx and P-Tx are significant sources of non-trivial control flows. Although external transactions account for only $12.18\%$ of the total, they introduce $52.78\%$ of non-trivial control flows, often bringing unexpected interactions and potential vulnerabilities. In contrast, privileged transactions, despite comprising $6.81\%$ of the total, contribute to $19.7\%$ of non-trivial flows, underscoring the critical role of administrative actions in protocol dynamics.
Answer to RQ3: CrossGuard effectively mitigates hacks by regulating non-trivial control flows, particularly those from external sources, demonstrating its critical role in enhancing DeFi security.
6.4. RQ4 and RQ5: Bypassability and Gas Overheads of CrossGuard
@@
Case Studies. Given that CrossGuard operates as a fully on-chain runtime system, it is transparent, allowing attackers to study its implementations and whitelisted control flows. A prevalent concern is whether informed attackers could bypass CrossGuard by splitting complex hack transactions into simpler ones. To address this, we perform in-depth studies of the $28$ hacks blocked by CrossGuard. Our analysis involves scrutinizing the control flows, underlying vulnerabilities, and the potential outcomes if attackers were to split their transactions.
@@
Results for Case Studies. The RQ5 column in Table 3 summarizes our findings: $25$ out of $28$ hacks cannot be bypassed by attackers. Specifically, $13$ hacks inherently require complex control flows to exploit vulnerabilities (marked as ✓).
@@
Additionally, 12 hacks rely on executing multiple capital-intensive functions to carry out the exploit. Historically, these attacks have used flash loans, which require all steps to be completed within a single transaction. Without flash loans, the attackers have to risk significant capital while competing with arbitrage bots, a scenario we deem as non-bypassable due to the high financial risks (marked as $\ast$). Only three hacks—Punk, DoughFina, and PikeFinance (marked as ✗)—show a bypass chance for attackers. The root causes of these exploits are access control, missing input validation, and re-initialization, respectively. Each of these vulnerabilities can be exploited by invoking a single function without requiring complex control flows. These attacks were initially caught by CrossGuard because the hackers included additional preparatory operations in their transactions. However, these preparatory steps can indeed be split and executed separately in separate transactions, allowing attackers to bypass CrossGuard.
Answer to RQ4: CrossGuard’s defenses remain effective against 25 out of 28 hacks.
@@
Experiment. To measure the gas overhead of CrossGuard, we instrumented smart contracts in a template-based manner. We insert specific code snippets at key points—such as function entry and exit, as well as during storage access operations. These snippets execute at runtime to capture the additional gas consumption introduced by CrossGuard. We used Foundry (Foundry Contributors, 2023) to compare the gas costs between the original and instrumented contracts, recording overhead differences in various execution phases. This process involved detailed assessments of each instrumentation type, functions within the guard contract, and EOA checks. During transaction replays, we logged the additional gas costs incurred at these key execution points to compute the overall gas overhead.
@@
Results for Gas Overheads. The RQ6 column in Table 3 presents the gas overhead introduced by CrossGuard for each benchmark. On average, the overall gas overhead is $15.52\%$. Notably, $14$ protocols exhibit a gas overhead below $5\%$, primarily due to the high proportion of EOA transactions within these benchmarks. Since EOA transactions benefit from Optimization 1 (Section 5.5), which reduces unnecessary gas consumption, the resulting gas overhead remains minimal. Even for the three widely used DeFi protocols, which feature a significant number of contract-initiated transactions, the average gas overhead remains 29.81%. This is a reasonable tradeoff considering the strong security guarantees provided by CrossGuard. The higher overhead in these cases stems from the need to track function calls and storage accesses, which are essential for securing protocols with complex execution flows.
Answer to RQ5: CrossGuard demonstrates a manageable gas overhead of 15.52% on average for the hacks benchmark.
6.5. RQ6: Comparative Analysis with the SOTA Tool Trace2Inv
@@
# Hacks Blocked 28 28 27 22
Avg. FP1.74 0.23 3.39 0.25
Results. Table 4 presents the analysis results that demonstrate that CrossGuard, even without training data, effectively blocks $28$ out of $30$ while maintaining an average FPrate of $1.74$%. This is a significant achievement compared to Trace2Inv, which requires training data to function but only blocks at most $27$ hacks. When trained, CrossGuard maintains its effectiveness in blocking hacks, reducing its FPrate dramatically to $0.23$%, which is superior to the FPrates achieved by Trace2Inv’s invariants ($3.39$% and $0.25$%). These results underline CrossGuard’s potential as a state-of-the-art solution providing robust security for DeFi applications. Moreover, CrossGuard and Trace2Inv can be used in conjunction to provide a more comprehensive security solution.
Answer to RQ6: CrossGuard surpasses Trace2Inv in hack prevention and false positive rates.
7. Threats to Validity
The internal threat to validity concerns potential human errors in identifying protected contracts. As discussed in Section 6.1, we rely on EtherScan labels to identify these core contracts to protect. The labels might be incomplete, leading to missing protected contracts. However, with more protected contracts identified, our approach will remain effective in blocking hack transactions, as the control flow of hack transactions will be more complex but still unique as per our approach. Our results may also face external threats due to the reliance on Trace2Inv (Chen et al., 2024) benchmarks, which focus on hacks up to June 2022. We mitigate this threat by including additional 8 hacks from February to July 2024. Additionally, we also include three major DeFi protocols—AAVE, Lido, and Uniswap—representing the most current protocols and user transactions.
@@ 8. Related Works
Invariant Generation and Enforcement. A significant body of research has focused on generating and enforcing invariants for smart contracts to ensure their security. Cider (Liu et al., 2022) employs deep reinforcement learning to derive invariants that prevent arithmetic overflows from contract source code. InvCon (Liu and Li, 2022) and InvCon+ (Liu et al., 2024) combine dynamic inference with static verification to produce verified, expressive contract invariants based on function preconditions and postconditions. Furthermore, Trace2Inv (Chen et al., 2024) dynamically learns invariants from transaction history, analyzing the effectiveness of $23$ invariants in preventing attacks. These studies predominantly focus on invariants tailored to individual contracts or specific kinds of vulnerabilities. In contrast, CrossGuard focuses on control flows across multiple contracts within the protocol. Unlike these prior studies, CrossGuard can achieve better true positive and false positive rates with a one-time configuration pre-deployment.
Control Flow Restriction. In the industry, SphereX (Spherex, 2024), a smart contract security firm, offers services to manually restrict control flows within smart contracts. While their goals align closely with ours, their approach requires protocol developers to manually select which control flows to whitelist or blacklist. This manual intervention contrasts sharply with our methodology, which not only automates the whitelisting of unseen control flows but also simplifies the overall control flow structure, significantly reducing the burden on developers and increasing the system’s adaptability.
Re-entrancy Attack Defense. Many researchers have proposed approaches to restrict smart contract control flows specifically to combat re-entrancy attacks. Tools like Oyente (Luu et al., 2016), Osiris (Torres et al., 2018), Reguard (Liu et al., 2018), Slither (Feist et al., 2019), MPro (Zhang et al., 2019), Sailfish (Bose et al., 2022), Pluto (Ma et al., 2021), Park (Zheng et al., 2022), SliSE (Wang et al., 2024), Albert et al., employ static analysis and symbolic execution to identify potential re-entrancy vulnerabilities, then apply re-entrancy guards as preventive measures. Notably, Sereum (Rodler et al., 2018) and Grossman et al. offer runtime validation frameworks to protect deployed contracts against re-entrancy attacks. Callens et al. have designed strategies to prevent the same function from being called twice within a transaction. Unlike these approaches, our work adopts a broader perspective on control flow restriction, targeting a more comprehensive set of vulnerabilities beyond just re-entrancy.
Secure Type System. Cecchetti et al. introduce a security type system that, alongside runtime mechanisms, robustly enforces secure information flow and re-entrancy controls, even amidst unknown code (Cecchetti et al., 2020, 2021). In contrast, our system is designed purely for runtime operation built entirely on the EVM, operating without the need for any supplementary type system, thereby simplifying integration and adoption.
@@ 9. Conclusion
In this paper, we presented CrossGuard, a novel control flow integrity framework specifically designed to secure smart contracts within the DeFi ecosystem. By instrumenting smart contract source code, CrossGuard dynamically prevents malicious transactions from executing risky control flow paths at runtime, effectively mitigating a broad spectrum of sophisticated attacks. Our comprehensive evaluation demonstrates that CrossGuard blocks the vast majority of benchmark attacks, significantly reduces false positives without relying on a pre-collected training set of benign transactions, and maintains a manageable gas overhead. Furthermore, integrating manual feedback enhances its accuracy, ensuring adaptability to emerging threats. Together, these results establish CrossGuard as a practical and robust solution for enhancing smart contract security, while providing valuable insights for the design of future DeFi protocols.

### v2

@@ 6.3. RQ3: Control Flows Introduced by Different Actors and False Positives
A key insight from this analysis is that a significant number of protocols ($27$, as highlighted in gray in the Total column) exhibit a relatively low number of non-trivial control flows ($\leq 9$) throughout their operational lifetimes prior to being hacked. Notably, in $17$ protocols (Warp, CheeseBank, XCarnival, Harvest, ValueDeFi, VisorFi, Eminence, IndexFi, RevestFi, Punk, DoughFina, BlueberryProtocol, PikeFinance, GFOX, OmniNFT, Auctus, MonoXFi), the hack was the first external non-trivial control flow introduced, beyond those generated internally (as indicated by gray cells showing 0 nCF from O-Tx but exactly $1$ nCF from E-Tx)  [^8] Two exceptions are XCarnival and Harvest. They were hacked in multiple hack transactions which introduced 3 and 2 non-trivial control flows, respectively.. This insight suggests that $17$ hacks could be simply prevented with zero false positive rates by restricting ALL external (“non-trusted”) developers, allowing only the protocol deployers to create new control flows. The analysis reveals that E-Tx and P-Tx are significant sources of non-trivial control flows. Although external transactions account for only $13.17\%$ of the total, they introduce $65.64\%$ of non-trivial control flows, often bringing unexpected interactions and potential vulnerabilities. In contrast, S-Tx, despite comprising $70.14\%$ of the total, contribute to only $20.01\%$ of non-trivial flows, underscoring the fact that most users interact directly with the protocol rather than through other intermediate contracts.
We also utilize the above identified transaction categories to conduct a deeper analysis of false positives in RQ2. We examined the three protocols with the highest false positive rates: bZx2 (3.57%), UwULend (2.38%), and RariCapital1 (1.22%)—all of which are lending protocols. Our investigation reveals that false positives predominantly originate from two specific actor categories: Other DeFi Protocols (O-Tx) accounting for 827 out of 1025 false positives in bZx2, 3 out of 468 in UwULend, and 74 out of 87 in RariCapital1; and External Actors (E-Tx) responsible for 198 out of 1025 in bZx2, 465 out of 468 in UwULend, and 13 out of 87 in RariCapital1. Additionally, we identified that 9 out of 13 false positives in RariCapital1 stem from MEV operations.
Further investigation into the nature of these false positives reveals two primary categories of legitimate use cases that introduce novel control flows. The first category involves helper contracts designed to streamline user operations, such as a contract that facilitates depositing three different tokens into UwULend by sequentially invoking the deposit function three times within a single transaction. The second category encompasses innovative functionality extensions, exemplified by patterns observed in UwULend where users deploy contracts that invoke deposit and borrow functions multiple times in sophisticated sequences to maximize their borrowing power and optimize capital efficiency. These legitimate but complex interaction patterns highlight opportunities for future refinement of control flow policies to better accommodate common DeFi usage patterns while maintaining security guarantees.
6.4. RQ4 & 5: Bypassability and Gas Overheads
@@
Case Studies. Given that CrossGuard operates as a fully on-chain runtime system, it is transparent, allowing attackers to study its implementations and whitelisted control flows. A prevalent concern is whether informed attackers could bypass CrossGuard by splitting complex hack transactions into simpler ones. To address this, we perform in-depth studies of the $35$ hacks blocked by CrossGuard. Our analysis involves scrutinizing the control flows, underlying vulnerabilities, and the potential outcomes if attackers were to split their transactions.
@@
Results for Case Studies. The RQ4 column in Table 3 summarizes our findings: $31$ out of $35$ hacks cannot be bypassed by attackers. Specifically, $18$ hacks inherently require complex control flows to exploit vulnerabilities (marked as ✓).
@@
Additionally, $13$ hacks rely on executing multiple capital-intensive functions to carry out the exploit. Historically, these attacks have used flash loans, which require all steps to be completed within a single transaction. Without flash loans, the attackers have to risk their own capital while competing with arbitrage bots, a scenario we deem as non-bypassable due to the high financial risks (marked as $\ast$). Only five hacks Punk, DoughFina, PikeFinance, Audius, and Auctus (marked as ✗) show a bypass chance for attackers. The root causes of these exploits are access control, missing input validation, and initialization bugs. Each of these vulnerabilities can be exploited by invoking a single function without requiring complex control flows. These attacks were initially caught by CrossGuard because the hackers included additional preparatory operations in their transactions. However, these preparatory steps can indeed be split and executed separately in separate transactions, allowing attackers to bypass CrossGuard.
@@
Experiment. To measure the gas overhead of CrossGuard, we instrumented smart contracts in a template-based manner. We insert specific code snippets at key points such as function entry and exit, as well as during storage access operations. These snippets execute at runtime to capture the additional gas consumption introduced by CrossGuard. We used Foundry (Foundry Contributors, 2023) to compare the gas costs between the original and instrumented contracts, recording overhead differences in various execution phases. This process involved detailed assessments of each instrumentation type, functions within the guard contract, and EOA checks. During transaction replays, we logged the additional gas costs incurred at these key execution points to compute the overall gas overhead.
To benchmark against industry standards, we evaluated the Hyperithm protocol (39; Etherscan (2025c); Etherscan (2025a); Etherscan (2025b)) deployed on May 22, 2025, which is protected by SphereX, an industry control flow restriction solution discussed in Section 8. We measured the additional gas costs during transaction replays to compute overhead for both CrossGuard and SphereX implementations on identical transaction history.
@@
Results for Gas Overheads. The RQ5 column in Table 3 presents the gas overhead introduced by CrossGuard for each benchmark. On average, the overall gas overhead is $3.53\%$. Notably, $28$ protocols exhibit a gas overhead below $5\%$, primarily due to the high proportion of EOA transactions within these benchmarks. Since EOA transactions benefit from Optimization 1 (Section 5.5), which reduces unnecessary gas consumption, the resulting gas overhead remains minimal. Even for the three widely used DeFi protocols, which feature a significant number of contract-initiated transactions, the average gas overhead remains $7.52\%$. This is a reasonable tradeoff considering the strong security guarantees provided by CrossGuard. The higher overhead in these cases stems from the need to track function calls and storage accesses, which are essential for securing protocols with complex execution flows.
Our industry comparison reveals that SphereX introduces 6.09% overhead for the Hyperithm protocol, while CrossGuard achieves only 0.22% overhead on the same contracts. Analysis of 166 transactions shows that 127 are simple invocations optimized by Optimization 1, and 38 are deployer-originated transactions optimized by Optimization 4. In contrast, SphereX lacks these optimizations and triggers external calls for every function invocation, resulting in substantially higher gas overhead. This demonstrates CrossGuard’s superior efficiency through targeted optimization strategies.
6.5. RQ6: Comparative Analysis with Trace2Inv
@@
# Hacks Blocked 35 35 34 29
Avg. FP% 1.19 0.15 3.14 0.23
Results. Table 4 presents the analysis results that demonstrate that CrossGuard, even without training data, effectively blocks $35$ out of $37$ while maintaining an average FP% rate of 1.191.19%. This is a significant achievement compared to Trace2Inv, which no only requires training data to function but only blocks at most $34$ hacks. When trained, CrossGuard maintains its effectiveness in blocking hacks, reducing its FP% rate dramatically to 0.150.15%, which is superior to the FP% rates achieved by Trace2Inv’s invariants (3.143.14% and 0.230.23%). These results underline CrossGuard’s potential as a state-of-the-art solution providing robust security for DeFi applications. Moreover, CrossGuard and Trace2Inv can be used in conjunction to provide a more comprehensive security solution.
7. Discussion and Threats to Validity
Generalization. CrossGuard can generalize through two mechanisms. First, it can evolve and reduce false positives by learning from blocked benign transactions, enabling administrators to whitelist new control flow patterns. Second, it records all non-trivial control flows in the guard contract. If novel attacks bypass current conditions (Alg. 1, Line 27), developers can update the guard contract to add new detection logic and automatically revert such transactions.
Integrating CrossGuard Pre- and Post-deployment. For new protocols, CrossGuard’s integration involves three steps: (1) instrument contracts (Section 5), using Slither and Heimdall to omit unnecessary instrumentation (Section 5.5); (2) deploy a guard contract to enforce policies and heuristics; and (3) configure access controls for all components. For existing deployed upgradable protocols (Bodell III et al., 2023), CrossGuard can be retrofitted by upgrading implementation contracts to new instrumented versions, thereby enabling CrossGuard without disrupting existing protocol logic. Following deployment, CrossGuard operates autonomously. Administrators can further improve CrossGuard by whitelisting blocked legitimate control flows (false positives). This adaptive process ensures that CrossGuard evolves alongside protocol growth, maintaining strong security guarantees over time.
Threats to Validity. The internal threat to validity concerns potential human errors in identifying protected contracts. As discussed in Section 6.1, we rely on Etherscan labels to identify these core contracts to protect. The labels might be incomplete, leading to missing protected contracts. However, with more protected contracts identified, our approach will only become more effective in blocking hacks, as the control flows of hack transactions recorded by CrossGuard will be more complex but still unique. Our results may also face external threats due to the reliance on Trace2Inv (Chen et al., 2024) and Sting (Zhang et al., 2023) benchmarks, which focus on hacks up to 2022. We mitigate this threat by including additional 8 hacks from February to July 2024. Additionally, we also include three major DeFi protocols representing the most current protocols and user transactions.
@@ 8. Related Works
Invariant Generation and Enforcement. Prior research generates and enforces invariants for contract security. Cider (Liu et al., 2022) derives arithmetic overflow invariants via deep reinforcement learning, while InvCon (Liu and Li, 2022) and InvCon+ (Liu et al., 2024) combine dynamic inference with static verification for function-level invariants. Trace2Inv (Chen et al., 2024) learns invariants from transaction history to prevent attacks. Unlike these approaches, which focus on individual contracts or specific bugs, CrossGuard is a general-purpose framework that targets control flows across multiple protocol contracts with a one-time pre-deployment configuration.
Control Flow Restriction. SphereX (Spherex, 2024) in industry offers services to manually restrict control flows, requiring developers to explicitly whitelist or blacklist paths. In contrast, CrossGuard automates the whitelisting of unseen control flows and simplifies the overall structure, significantly reducing developer burden and increasing system adaptability compared to manual intervention.
Re-entrancy Attack Defense and Secure Type System. Numerous tools restrict control flows to combat re-entrancy attacks. Static analysis tools (Luu et al., 2016; Torres et al., 2018; Liu et al., 2018; Feist et al., 2019; Zhang et al., 2019; Bose et al., 2022; Ma et al., 2021; Zheng et al., 2022; Wang et al., 2024; Albert et al., 2020) identify re-entrancy vulnerabilities and apply guards. Runtime frameworks like Sereum (Rodler et al., 2018) and Grossman et al. protect deployed contracts, while Callens et al. prevent duplicate function calls within transactions. Unlike these re-entrancy-focused approaches, CrossGuard adopts broader control flow restriction targeting comprehensive vulnerability classes. Cecchetti et al. propose a security type system to enforce information flow and re-entrancy controls (Cecchetti et al., 2020, 2021). Conversely, CrossGuard is a purely runtime system built entirely on the EVM, operating without supplementary type systems or language modifications.
@@ 9. Conclusion
In this paper, we presented CrossGuard, a novel control flow integrity framework specifically designed to secure DeFi smart contracts at runtime. Configured only once at deployment, CrossGuard prevents malicious transactions from executing risky control flows on the fly, effectively mitigating a wide range of attacks. Our comprehensive evaluation demonstrates that CrossGuard blocks the vast majority of benchmark attacks, significantly reduces false positives without relying on a pre-collected training set of benign transactions, and maintains a manageable gas overhead. Furthermore, integrating manual feedback enhances its accuracy. Together, these results establish CrossGuard as a practical solution for securing smart contracts. The artifact is hosted at https://github.com/jeffchen006/CrossGuard-Artifact (DOI: 10.5281/zenodo.18262998).
Acknowledgements.
We thank anonymous reviewers for their insightful comments on the early version of the paper. This work was supported by Mitacs through the Mitacs Accelerate program.